OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12

hartnn (hartnn):

start by dividing -3 on both sides, what do u get ?

OpenStudy (anonymous):

4.

hartnn (hartnn):

you mean |2x+6| = 4 ?

OpenStudy (anonymous):

Right.

hartnn (hartnn):

now use this if |a| = b then a= b or a = -b

hartnn (hartnn):

so if |2x+6| = 4 2x+6 = 4 or 2x+6 = - 4 can you solve these ?

OpenStudy (anonymous):

|2x+6| = -12 or |2x+6|= 12

hartnn (hartnn):

when you split, you remove the absolute signs

hartnn (hartnn):

and it won't be 12 on right side, isn't it ? we divided by -3 and got 4 on right side...

OpenStudy (anonymous):

2x+6 = 4 x = -5

OpenStudy (anonymous):

So, x = 5 and x = -5 I just got that.

OpenStudy (anonymous):

I mean x = -1

hartnn (hartnn):

2x+6 = 4 gives x = -1 or 2x+6 = - 4 gives x = -5 so, yes, now its time to verify

hartnn (hartnn):

plug in x = -1 in your original equation on left side see if you get right side

OpenStudy (anonymous):

-3|2(-1)+6| = -12

OpenStudy (anonymous):

Yes I do.

hartnn (hartnn):

now try x = -5

OpenStudy (anonymous):

I got 12.

hartnn (hartnn):

how ? remember |-4| = 4

hartnn (hartnn):

because |a| = a if a is positive |a| = - a if a is negative

OpenStudy (anonymous):

I did -5 x 2 + * -3

OpenStudy (anonymous):

+6*

hartnn (hartnn):

you didn't consider the absolute sign ?

OpenStudy (anonymous):

I did everything in the absolute sign 1st.

OpenStudy (anonymous):

then mutiplid it by -3.

hartnn (hartnn):

|2(-5)+6|* (-3) = |-10 +6 | *(-3) = |-4|*(-3) = 4 *(-3) =-12

OpenStudy (anonymous):

Aren't we doing the full original equation?

OpenStudy (anonymous):

Ahhhhhh.

OpenStudy (anonymous):

Absolute number changes to a positive?

hartnn (hartnn):

so , you just verified that -1 and -5 are both the correct solutions to your equation :) any doubts ?

hartnn (hartnn):

yes, thats correct... like |-12| = 12 it keeps positive no. as positive |12| = 12

OpenStudy (anonymous):

Alright 1 more thing here.

OpenStudy (anonymous):

You can help me with it? Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. Equation 1 Equation 2 |5x + 6| = 41 |2x + 13| = 28 Equation 1 has more solutions than equation 2. Equation 1 and Equation 2 have the same number of solutions. Equation 2 has more solutions than Equation 1. The number of solutions cannot be determined.

hartnn (hartnn):

you solve both the equation, in the exact same way we solved −3|2x + 6| = −12

hartnn (hartnn):

then verify your solutions.

OpenStudy (anonymous):

They have the same number of solutions?

hartnn (hartnn):

thats correct! :)

OpenStudy (anonymous):

Awesome, thanks so much for the help.

hartnn (hartnn):

welcome ^_^